The following inventory information is available:

Transaction

Date

Units

Dollars

Beginning Balance

1/1/2019

800

$6.00

Purchase

1/10/2019

600

$5.50

Purchase

1/20/2019

950

$5.25

Purchase

1/25/2019

350

$5.00

Sale throughout month

1700

$11.75

Calculate cost of goods available for sale, cost of goods sold, gross profit, gross profit margin and ending inventory using average cost, LIFO and FIFO inventory methods. The company uses the periodic inventory system. Then state which inventory method would provide the highest net income and Which inventory method provides the highest net income?  

Average Cost

LIFO

FIFO

Cost of goods available for sale

Cost of goods sold

Gross profit

Gross profit margin

Ending inventory

  1. A piece of equipment with a net book value of $45,000 and an original cost of $63,000 was sold for a $2,500 gain. The equipment had a salvage value of $9,000 and an estimated useful life of 9 years. The company used straight line depreciation.
    1. What was the accumulated depreciated of this equipment when the equipment was sold?
    2. What was the equipment sold for?
    3. What is the annual straight line depreciation expense related to this equipment?
    4. How many years did the company own and depreciate the equipment?

Expert Answer
Answer #1
FIFO Cost of Goods Available for sale Cost of Goods Sold Ending Inventory
No of Units Cost per unit Cost of Goods Available for sale No of Units Cost per unit Cost of Goods Sold No of units Cost per unit Ending Inventory
Beginning Inventory 800 $              6.00 $              4,800.00 800 $                   6.00 $      4,800.00
Purchases :
1/10/19 600 $              5.50 $              3,300.00 600 $                   5.50 $      3,300.00
1/20/19 950 $              5.25 $              4,987.50 300 $                   5.25 $      1,575.00 650 $       5.25 $       3,412.50
1/25/19 350 $              5.00 $              1,750.00 350 $       5.00 $       1,750.00
Total 2700 $            14,837.50 1700 $      9,675.00 1000 $       5,162.50
LIFO Cost of Goods Available for sale Cost of Goods Sold Ending Inventory
No of Units Cost per unit Cost of Goods Available for sale No of Units Cost per unit Cost of Goods Sold No of units Cost per unit Ending Inventory
Beginning Inventory 800 $              6.00 $              4,800.00 800 $       6.00 $       4,800.00
Purchases :
1/10/19 600 $              5.50 $              3,300.00 400 $                   5.50 $      2,200.00 200 $       5.50 $       1,100.00
1/20/19 950 $              5.25 $              4,987.50 950 $                   5.25 $      4,987.50
1/25/19 350 $              5.00 $              1,750.00 350 $                   5.00 $      1,750.00
Total 2700 $            14,837.50 1700 $      8,937.50 1000 $       5,900.00

Average Cost per unit = $14837.50 / 2700 = $5.50 per unit

Average Cost LIFO FIFO
Cost of goods available for sale $        14,837.50 $    14,837.50 $            14,837.50
Cost of goods sold $          9,342.13 $      9,675.00 $              8,937.50
Gross profit $          5,495.37 $      5,162.50 $              5,900.00
Gross profit margin 37.04% 34.79% 39.76%
Ending inventory $          5,495.37 $      5,162.50 $              5,900.00

1.
a. Accumulated Depreciation = Cost of asset - Net book value
= $63000 - 45000 = $18000

b. Equipment sale proceeds = Net Book Value + Gain
= $45000 + 2500 = $47500

c. Straight line Depreciation = (Original value - Salvage Value) / Useful life
= ($63000-9000)/9 = $6000

d. Years company own asset = $18000 /6000 = 3 years